please try to solve for all, but if you don't have time try for b for sure. :D
(0.01a+0.05b+0.10c+0.25d)/(a+b+c+d)=(0.01(a-1)+0.05b+0.10c+0.25d)/((a-1)+b+c+d)

Answer by venugopalramana(3286)

please try to solve for all, but if you don't have time try for b for sure. :D
(0.01a+0.05b+0.10c+0.25d)/(a+b+c+d)=(0.01(a-1)+0.05b+0.10c+0.25d)/((a-1)+b+c+d)
CROSS MULTIPLYING..
[0.01a+0.05b+0.10c+0.25d)(a+b+c+d-1)]=[0.01(a-1)+0.05b+0.10c+0.25d][a+b+c+d)
[0.01a+0.05b+0.10c+0.25d][a+b+c+d]-[0.01a+0.05b+0.10c+0.25d]=
=[0.01a+0.05b+0.10c+0.25d)[a+b+c+d]-[a+b+c+d)
0.01a+0.5b+0.1c+0.25d=a+b+c+d
0.99a+0.5b+0.9c+0.75d=0
hence you can have any value for 3 of the 4 variables a,b,c and d....,the 4th.variable,dependent on the other 3...say
you can have a=b=c=0 say
then d=0
if a=0=b,c=1
d=-0.9/0.75=-0.3/0.25 =-1.2..and in general
d=-(0.99a+0.5b+0.9c)/0.75,for any value of a,b, and c will be a solution set.